Bait-N-Tackle sells fishing equipment. One of the company’s products, a basic tackle box, sells for $48 per unit. Variable expenses are $36 per tackle box, and fixed expenses associated with the tackle box total $18,000 per month.

Required:

1. Compute the company’s break-even point in number of tackle boxes and in total sales dollars.

2. If the variable expenses per tackle box increase as a percentage of the selling price, will it result in a higher or a lower break-even point? (Assume that the fixed expenses remain unchanged.)

  • Higher break-even point

  • Lower break-even point

3. At present, the company is selling 2,600 tackle boxes per month. The sales manager is convinced that a 12.5% reduction in the selling price will result in a 20% increase in the number of tackle boxes sold each month. Prepare two contribution income statements, one under present operating conditions, and one as operations would appear after the proposed changes.

4. Refer to the data in requirement 3 above. How many tackle boxes would have to be sold at the new selling price to yield a minimum pre-tax operating income of $14,400 per month?

5. Refer to the data in requirement 3 above. How many tackle boxes would have to be sold at the new selling price to generate after-tax operating income of $16,800 if the tax rate is 30%?
Answer #1

1.) Breakeven point :-

Boxes = Fixed cost/contribution per unit

= 18,000/(48 - 36)

= 1,500 boxes per month

Dollar sales = Fixed cost/contribution margin ratio

= 18,000/(12/48 x 100)

= $72,000 per month

2.) Higher break even point

3.) Contribution margin statement :-

Present conditions
Sales (2,600 x 48) 124,800
Variable cost (2,600 x 36) 93,600
Contribution margin 31,200
Fixed cost 18,000
Net Income 13,200
Proposed conditions
Sales (3,120 x 42) 131,040
Variable cost (3,120 x 36) 112,320
Contribution margin 18,720
Fixed cost 18,000
Net Income 720

Proposed selling price = 48 x (1 - 0.125) = 42

Proposed sale unit = 2,600 x 1.20 = 3,120

4.) Contribution margin per unit in new selling price = 42 - 36 = $6

Units need to be sold

= (required profit before tax + fixed costs)/contribution per unit

= (14,400 + 18,000)/6

= 5,400 boxes

5.) Before tax profit = 16,800/(1 - 0.30) = 24,000

Units need to be sold = (24,000 + 18,000)/6

= 7,000 boxes